# G-Transformer: Counterfactual Outcome Prediction under Dynamic and Time-varying Treatment Regimes

### v2

@@ 3.5 Evaluation
We applied our proposed G-Transformer approach to predicting outcomes of sepsis patients in the ICU under alternative fluid resuscitation treatment regimes (e.g. aggressive vs. conservative) using de-identified real-world intensive care units (ICU) data from the MIMIC-IV (Johnson et al., 2023b) publicly available from PhysioNet.
We quantitatively evaluate G-Transformer’s performance in predicting patient trajectories under observational treatment regimes learned from the ICU data. This entails using the G-Transformer architecture to predict treatment at each time-step under the ”usual care” in the observational data. Under predictive check, the treatments are from G-Transformer predicted actions conditioned on patient history.
@@ 4.1 Cancer Growth Experiments and Results
As in Lim et al. (2018); Bica et al. (2020a); Li et al. (2021); Melnychuk et al. (2022), we generate simulated ‘observational’ data from a pharmacokinetic-pharmacodynamic model of tumor growth under a stochastic regime (Geng et al., 2017). In this simulation, chemotherapy and radiation therapy comprise a two dimensional time-varying treatment impacting tumor growth. Under the observational regime, probability of receiving each treatment at each time depends on volume history, so there is time-varying confounding. We use the same experiment settings and data generation procedure as reported in Li et al. (2021). We generated four test sets in which four counterfactual regimes were followed for the final four time points in the test set: (1) give only radiotherapy, (2) only chemotherapy, (3) both chemotherapy and radiotherapy, and (4) no treatment. These counterfactual regimes are static time-varying regimes (in which treatments do not depend on prior patient history), as rMSN, CRN, and Causal Transformer are intended to estimate counterfactuals under static time-varying regimes.
Table 1 presents the percent root mean square error (RMSE) of predictions from rMSN, CRN, linear implementation of g-computation, G-Net, Causal Transformer, and G-Transformer in the final four time steps when counterfactual strategies were in effect, conditioned upon previous time points. The raw RMSE values were divided by the maximum possible tumor volume, 1150 $cm^{2}$, as in Lim et al. (2018), to calculate the percent RMSE.
We observe that G-Transformer achieved the best overall RMSE performance in three of the four cancer growth datasets, demonstrating its advantages in counterfactual prediction task compared to other state-of-the-art deep learning models, including rMSN, CRN, G-Net and Causal Transformer. Causal Transformer outperformed other models under the chemotherapy (Chemo) counterfactual regime, but fall short in all other counterfactual regimes.
@@
rMSN CRN Linear G-Net CT Ours
$t$ g-comp G-Transformer
@@ 4.2 CVSim Experiments and Results
Given observed covariate history through 34 time steps and treatment history through 33 time steps of each trajectory in each $D_{cj}$, we computed both the population-level and individual-level RMSE, as well as the calibration, of counterfactual predictions from G-Transformer and its counterparts for time steps 35 to 66. Moreover, we presented the estimated and actual population-level average trajectories under gc1 and gc2 for selected variables. We also visually demonstrated the treatment effects estimation.
Population-Level and Individual-Level RMSE
We use both population-level and individual-level RMSE to assess the predictive accuracy of the model for continuous variables in CVSim data. To derive population-level RMSE, we aggregated the simulated trajectories from all patients into a single population-level trajectory. This predicted average population-level trajectory was then compared to the ground-truth average population-level trajectory for RMSE calculation. Individual-level RMSE is calculated by first determining the square root of the mean squared differences between each patient’s observed and predicted trajectories. For G-Transformer, G-Net and the linear implementation of g-computation, the predicted trajectory is averaged from 100 Monte Carlo simulations conditioned on historical data up to time (k-1). The RMSE values across all patients are then averaged to derive individual-level RMSE.
Table 2 presents a comparison of individual-level and population-level RMSE among linear implementation of g-computation, G-Net, Causal Transformer, and G-Transformer under gc1 and gc2. The G-Transformer outperforms linear implementation of g-computation and G-Net, and considerably outperforms Causal Transformer in this counterfactual prediction task. It should be noted that Causal Transformer Melnychuk et al. (2022) can estimate counterfactual outcomes under time-varying treatments, but it is not designed to estimate counterfactual outcomes under dynamic treatment regimes, where treatments depend on past covariate history. Nevertheless, we present its performance here as a baseline.
Table 2: CVSim: Counterfactual prediction under gc1 and gc2. Performance of various models in predicting remaining 32-hour patient covariate trajectories based on first 34-hour patient covariate trajectories. Values reported represent the population-level and individual-level RMSE between the predicted and counterfactual trajectories.
@@
Linear(g-comp) 0.252 0.272 1.103 1.193
G-Net 0.061 0.105 1.020 1.114
Causal Transformer 0.885 1.099 1.435 1.727
G-Transformer 0.050 0.054 1.016 1.104
Calibration
Figure 2 depicts calibration for G-Transformer, G-Net, and linear implementation of g-computation with lower and upper quantiles of 0.05 and 0.95. G-Transformer generally performed better than G-Net and linear implementation of g-computation over all time steps. The model is well calibrated at the beginning, but as simulation time steps increases, calibration becomes weaker for all models. This is in line with the RMSE increase over time as can be seen from Table 7 in the Appendix A.3.
(a) Calibration under $gc_{1}$
(b) Calibration under $gc_{2}$
Figure 2: Calibration over time for $g_{c1}$ and $g_{c2}$
@@
The G-Transformer, G-Net, and linear implementation of g-computation can be used to quantitatively demonstrate counterfactual outcomes through population average trajectories. The estimated and actual population-level average trajectories under gc1 and gc2 for selected variables, specifically arterial flow (AQ), arterial pressure (AP), and heart rate (HR), are presented in Figure 3 (3(a)) to (3(f)). The plots show that G-Transformer outperforms G-Net and linear implementation of g-computation, and more accurately predict AQ, AP, and HR counterfactual trajectories from the population level.
@@
(g) AQ ($gc_{1}$-$gc_{2}$)
(h) AP ($gc_{1}$-$gc_{2}$)
(i) HR ($gc_{1}$-$gc_{2}$)
Figure 3: Plots (3(a)) to (3(f)) show the estimated and actual population average trajectories for three selected variables: arterial flow (AQ), arterial pressure (AP), and heart rate (HR), according to three models’ predictions under gc1 vs gc2. Plots (3(g)) to (3(i)) show population-level treatment effect of $g_{c1}$ vs $g_{c2}$ estimated by G-Transformer vs G-Net.
Treatment Effects
The G-Transformer, G-Net, and linear g-computation implementation can also be used to estimate population average counterfactual outcomes and treatment effects, quantities sometimes more relevant to policy decisions than individual-level counterfactual predictions. Figure 3 (3(g)) to (3(i)) show the estimates and true values of the population average treatment effect of following $g_{c1}$ as opposed to $g_{c2}$ on selected variables. We see that the G-Transformer effectively estimates these population-level quantities of interest.
@@ 5.2 Study Design
Since there is no ground-truth observations available for the MIMIC data for counterfactual predictions, we first assess our models’ performance through a “predictive check.” This involved conducting Monte Carlo simulations and projecting forward patient covariate trajectories within the test set under the observational regime. Subsequently, we compared these simulated trajectories (averaged across 100 Monte Carlo simulations per patient) to the ground-truth data. Following predictive check, we assess our model in performing counterfactual predictions.
Predictive Checks
@@
We compared the 24-hour trajectories between predicted and ground-truth under G-Transformer, G-Net, and Causal Transformer, using individual-level and population-level RMSE as metrics. The method for calculating the population-level RMSE is as described in the previous section. However, since patient trajectories can be censored due to events such as death and release in MIMIC data, individual-level RMSE is adjusted to account for error due to over- and under-prediction in the trajectory length as in Su et al. (see Appendix B.1 for details).
@@
We also compared the models’ performance in predicting the occurrence of adverse clinical outcomes during the first 24-hours patients were in the ICU. Confidence intervals around AUCs are generated based on techniques described in DeLong et al. (1988). To generate a probability for the 24-hour window, we first assigned a binary label for each Monte Carlo simulation based on whether the outcome is predicted to have occurred: $1$ if outcome of interest was predicted to occur between timestamps $k$ to $24$, and 00 otherwise. Then the average across the Monte Carlo simulations per patient was used as the predicted probability for that patient. The ground truth was determined in the same manner by looking for the presence of the outcome of interest in the ground-truth trajectory.
Counterfactual Experiments
We adapted our counterfactual strategies from established clinical trials studying the early treatment of sepsis. This strategy imposes a fluid cap on the total amount of fluids intake of a patient to $X$ liters, and ceases administration once the fluid cap is reached or when the patient is fluid-overloaded. More specifically, this strategy was based on the Crystalloid Liberal or Vasopressors Early Resuscitation in Sepsis (CLOVERS) clinical trial (Self et al., ). For a patient with blood pressure below 65mmHg at time $t$, a 1000mL bolus was administered if the total volume of fluids (including both treatment and maintenance) they received up until that time point did not exceed $X$ liters and if they did not exhibit any signs of fluid overload (as indicated by the presence of pulmonary edema based on chest x-ray radiology reports). We experimented with imposing a fluid cap of $3$L and $5$L as fluid conservative strategies, and a fluid strategy with no fluid cap to represent a fluid liberal strategy.
5.3 Predictive Check: Population-Level and Individual-Level RMSE
@@
Table 3 presents the predictive check on population-level and individual-level RMSE for continuous variables. We aim to assess the performance of G-Transformer, G-Net, and Causal Transformer in predicting 24-hour patient covariate trajectories, starting at time $k$ conditioned on covariates from the previous $k-1$ timesteps of post-ICU admission. $k$ ranges from 2 to 8 hours after ICU admission. Values reported represent the RMSE of continuous covariates between the predicted and ground-truth trajectories. To account for mismatch in trajectory length due to over- and under-prediction of death and release time, for both predicted and actual trajectories, depending on nature of covariates, post-death timesteps are padded with one of the following values: normalized zero, normalized minimum value in dataset, or normalized logarithmic zero. The post-release timesteps are consistently padded with normalized population mean.
@@
According to Table 3, we can see that errors in predicted 24-hour patient trajectories are relatively stable for G-Transformer and Causal Transformer, and G-Transformer consistently outperforms G-Net and Causal Transformer across different values of $k$. Although Causal Transformer Melnychuk et al. (2022) is not intended to estimate counterfactual outcomes under dynamic treatment regimes (where treatments depend on past covariate history), nevertheless, we present its performance here as a baseline.
@@
Table 3: MIMIC-IV Experiments: Predictive check on continuous variables at population-level and individual-level RMSE. $k$ is the simulation start time relative to the ICU admission time. Performance in predicting 24-hour patient covariate trajectories, starting at hour $k$ conditioned on covariates from the previous $k-1$ timesteps of post-ICU admission. $k$ ranges from 2 to 8 hours after ICU admission. Values reported represent the population-level and individual-level RMSE of continuous covariates between the predicted and ground-truth trajectories. Individual-level RMSE is adjusted to account for error due to over- and under-prediction in the trajectory length as a result of death and discharge.
Population Level RMSE Individual Level RMSE
Model / k (hours) 2 4 6 8 2 4 6 8
G-Net 0.338 0.265 0.214 0.181 15.85 13.71 11.40 9.29
Causal Transformer 2.770 2.736 2.643 2.514 14.37 15.00 15.73 16.54
G-Transformer 0.297 0.233 0.204 0.178 7.96 8.83 9.46 8.62
@@
Figure 4: Illustration of predictive check on an example MIMIC test set patient. G-Transformer simulated systolic blood pressure (SBP) and heart rate (HR) trajectories (100 Monte Carlo simulations in light blue, average in solid dark blue) compared with ground truth (red) for one patient under predictive check (simulation starting at hour 6).
@@
Table 4 presents AUCs (and 95% CI) in predicting clinical outcomes within the first 24 hours of ICU admission at varying simulation start times from G-Transformer and G-Net. There are six outcomes of interest: diagnosis of pulmonary edema, use of a mechanical ventilator (MV), administration of diuretics, dialysis, ICU release, and in-hospital mortality. We observe that the predicted 24-hour clinical outcomes of interest tend to be more accurate as the value of $k$ increases, and G-Transformer generally outperforms G-Net across different values of $k$.
@@
Table 4: Predictive Check. Performance of G-Transformer and G-Net in predicting clinical outcomes within the first 24 hours of ICU admission at varying simulation start times. Values reported are AUCs and 95% confidence intervals.
Model k Edema MV Diuretics Dialysis Release Death
G-Net 2 0.82 (0.79, 0.86) 0.89 (0.87, 0.91) 0.68 (0.63, 0.73) 0.85 (0.75, 0.95) 0.67 (0.61, 0.73) 0.79 (0.67, 0.90)
G-Transformer 2 0.84 (0.80, 0.87) 0.89 (0.86, 0.91) 0.73 (0.68, 0.78) 0.90 (0.81, 0.98) 0.67 (0.61, 0.73) 0.82 (0.71, 0.93)
G-Net 4 0.86 (0.83, 0.89) 0.94 (0.92, 0.95) 0.70 (0.64, 0.75) 0.94 (0.88, 1.00) 0.65 (0.59, 0.71) 0.86 (0.76, 0.96)
G-Transformer 4 0.86 (0.83, 0.89) 0.94 (0.92, 0.96) 0.75 (0.70, 0.80) 0.96 (0.91, 1.00) 0.68 (0.62, 0.74) 0.91 (0.83, 1.00)
G-Net 6 0.89 (0.86, 0.92) 0.95 (0.93, 0.96) 0.73 (0.67, 0.78) 0.93 (0.86, 1.00) 0.68 (0.63, 0.74) 0.83 (0.72, 0.94)
G-Transformer 6 0.87 (0.84, 0.90) 0.94 (0.92, 0.96) 0.74 (0.69, 0.80) 0.96 (0.90, 1.00) 0.69 (0.63, 0.74) 0.89 (0.80, 0.98)
G-Net 8 0.89 (0.87, 0.92) 0.95 (0.94, 0.97) 0.73 (0.68, 0.79) 0.97 (0.92, 1.00) 0.71 (0.66, 0.77) 0.89 (0.79, 0.98)
G-Transformer 8 0.89 (0.86, 0.91) 0.95 (0.93, 0.97) 0.75 (0.70, 0.81) 0.95 (0.89, 1.00) 0.71 (0.66, 0.77) 0.95 (0.79, 0.98)
Counterfactual Prediction: Population-Level Trajectories

### v4

@@ 3.5 Evaluation
We applied our proposed G-Transformer approach to predicting outcomes of sepsis patients in the ICU under alternative fluid resuscitation treatment regimes (e.g. aggressive vs. conservative) using de-identified real-world intensive care units (ICU) data from the MIMIC-IV (Johnson et al., 2023b) publicly available from PhysioNet Goldberger et al. (2000). We quantitatively evaluate G-Transformer’s performance in predicting patient trajectories under observational treatment regimes learned from the ICU data. This entails using the G-Transformer architecture to predict treatment at each time-step under the observational treatment regime in the observational data. Under predictive check, the treatments are from G-Transformer predicted actions conditioned on patient history, and we evaluate by comparing the predicted time-varying outcomes (averaged across $M$ Monte-Carlo simulations) with the actual observed trajectories from individual patients.
@@ 4.1 Cancer Growth Experiments and Results
As in Lim et al. (2018); Bica et al. (2020a); Li et al. (2021); Melnychuk et al. (2022), we generate simulated ‘observational’ data from a pharmacokinetic-pharmacodynamic (PK/PD) model of tumor growth under a stochastic regime (Geng et al., 2017). In this simulation, chemotherapy and radiation therapy comprise a two dimensional time-varying treatment impacting tumor growth. Under the observational regime, probability of receiving each treatment at each time depends on volume history, so there is time-varying confounding.
@@
rMSN CRN Linear G-Net CT G-Transformer
$t$ g-comp
@@
We use the same experiment settings and data generation procedure as reported in Li et al. (2021). Briefly, we generated four test sets in which four counterfactual regimes were followed for the final four time points in the test set: (1) give only radiotherapy, (2) only chemotherapy, (3) both chemotherapy and radiotherapy, and (4) no treatment. We generated 10,000 trajectories for training, and 1,000 simulated trajectories for hyperparameter optimization (validation data), with another 1,000 for testing. See Li et al. (2021) for full details of the data generation process. These counterfactual regimes are static time-varying regimes (in which treatments do not depend on prior patient history), as rMSN, CRN, and Causal Transformer are intended to estimate counterfactuals under static time-varying regimes.
Table 1 presents the percent root mean square error (RMSE) of predictions from rMSN, CRN, linear implementation of g-computation, G-Net, Causal Transformer, and G-Transformer in the final four time steps when counterfactual strategies were in effect, conditioned upon previous time points. The raw RMSE values were divided by the maximum possible tumor volume, 1150 $cm^{2}$, as in Lim et al. (2018), to calculate the percent RMSE. Details of the model training and hyper-parameter settings are described in the Appendix.
We observe that G-Transformer achieved the best overall RMSE performance in three of the four cancer growth datasets, demonstrating its advantages in counterfactual prediction task compared to other state-of-the-art deep learning models, including rMSN, CRN, G-Net and Causal Transformer. Causal Transformer outperformed other models under the chemotherapy (Chemo) counterfactual regime, but fall short in other counterfactual regimes.
@@ 4.2 CVSim Experiments and Results
Given observed covariate history through 34 time steps and treatment history through 33 time steps of each trajectory in each $D_{cj}$, we computed both the population-level and individual-level RMSE, as well as the calibration, of counterfactual predictions from G-Transformer and its counterparts for time steps 35 to 66. Moreover, we presented the estimated and actual population-level average trajectories under gc1 and gc2 for selected variables.
CVSim Counterfactual Prediction Performance in RMSE.
We use RMSE to assess our model’s accuracy in counterfactual prediction for continuous variables in CVSim data. Individual-level RMSE is calculated by first determining the square root of the mean squared differences between each patient’s observed and predicted trajectories. We also report population-level RMSE, by comparing the average population-level trajectory of the predicted and the ground-truth.
For G-Transformer, G-Net and the linear implementation of g-computation, the predicted trajectory is averaged from 100 Monte Carlo simulations. The RMSE values across all patients are then averaged to derive individual-level RMSE.
Table 2: CVSim: Counterfactual prediction under gc1 and gc2. Performance of various models in predicting remaining 32-hour patient covariate trajectories based on first 34-hour patient covariate trajectories. Values reported represent the population-level and individual-level RMSE between the predicted and counterfactual trajectories. * Causal Transformer modified to update treatment assignment at each time-step under gc1 and gc2.
@@
Causal Transformer 0.576 $\pm$ 0.078 1.323 $\pm$ 0.257 1.473 $\pm$ 0.272 1.694 $\pm$ 0.407
Linear(g-comp) 0.255 $\pm$ 0.001 0.442 $\pm$ 0.002 1.103 $\pm$ 0.001 1.307 $\pm$ 0.001
G-Net 0.066 $\pm$ 0.001 0.329 $\pm$ 0.101 1.020 $\pm$ 0.002 1.220 $\pm$ 0.048
G-Transformer 0.055 $\pm$ 0.006 0.299 $\pm$0.110 1.019 $\pm$ 0.002 1.210 $\pm$ 0.045
Table 2 presents a comparison of individual-level and population-level RMSE among linear implementation of g-computation, G-Net, Causal Transformer, and G-Transformer under gc1 and gc2. The G-Transformer outperforms linear implementation of g-computation and G-Net, and considerably outperforms Causal Transformer in this counterfactual prediction task. Table 6 in the Appendix A.2 presents performance in RMSE over-time. It should be noted that Causal Transformer Melnychuk et al. (2022) can estimate counterfactual outcomes under time-varying treatments, but it is not designed to estimate counterfactual outcomes under dynamic treatment regimes, where treatments depend on past covariate history during counterfactual prediction. Nevertheless, we modified the Causal Transformer’s implementation to update the sequential treatment assignment at each time-step according to the specified dynamic treatment rule depending on the past covariates, and present its performance here as a baseline.
@@
Figure 2: Plots 2(a) to 2(f) show the estimated and actual population average trajectories for three selected variables: arterial flow (AQ), arterial pressure (AP), and heart rate (HR), according to three models’ predictions under gc1 vs gc2.
The G-Transformer, G-Net, and linear implementation of g-computation can be used to quantitatively demonstrate counterfactual outcomes through population average trajectories. The estimated and actual population-level average trajectories under gc1 and gc2 for selected variables, specifically arterial flow (AQ), arterial pressure (AP), and heart rate (HR), are presented in Figure 2 2(a) to 2(f). The plots show that G-Transformer outperforms G-Net and linear implementation of g-computation, and more accurately predict AQ, AP, and HR counterfactual trajectories from the population level.
Calibration
(a) Calibration under $gc_{1}$
(b) Calibration under $gc_{2}$
Figure 3: Calibration over time for $g_{c1}$ and $g_{c2}$
Figure 3 depicts calibration for G-Transformer, G-Net, and linear implementation of g-computation with lower and upper quantiles of 0.05 and 0.95. G-Transformer generally performed better than G-Net and linear implementation of g-computation over all time steps.
@@ 5.2 Study Design
Since there is no ground-truth observations available for the MIMIC data for counterfactual predictions, we first assess our models’ performance through a “predictive check” using data collected during the first 24 hours of patients’ stay in the ICU. This involved conducting Monte Carlo simulations and projecting forward patient covariate trajectories within the test set under the observational regime. Subsequently, we compared these simulated trajectories (averaged across 100 Monte Carlo simulations per patient) to the ground-truth data. Following predictive check, we assess our model in performing counterfactual predictions. In the MIMIC experiment, we use a separate Transformer encoder for each covariate in G-Transformer. For G-Net, we use a separate LSTM for each covariate.
5.3 Predictive Checks
@@
For each patient in the test set, we use G-Transformer and other baseline approaches to simulate their trajectories under the observational treatment regime from hour k up to hour 24, conditioned on the observed covariates and treatment history up to and including hour k-1. Under predictive check in G-Transformer, we use a Transformer encoder to predict the treatment action at each time step under the observational treatment regime. Similarly, for G-Net, we use an LSTM model to predict the treatment at each time step during predictive checks. We compared the trajectories between predicted and ground-truth under G-Transformer, G-Net, and Causal Transformer. We use individual-level RMSE as a metric. Since patient trajectories can be censored due to events such as death and release in MIMIC data, individual-level RMSE is adjusted to account for error due to over- and under-prediction in the trajectory length as in Su et al. (see Appendix B.1 for details).
@@
We also compared the models’ performance in predicting the occurrence of adverse clinical outcomes during the first 24-hours patients were in the ICU. To generate a probability for the 24-hour window, we first assigned a binary label for each Monte Carlo simulation based on whether the outcome is predicted to have occurred: $1$ if outcome of interest was predicted to occur between timestamps $k$ to $24$, and 00 otherwise. Then the average across the Monte Carlo simulations per patient was used as the predicted probability for that patient. The ground truth was determined in the same manner by looking for the presence of the outcome of interest in the ground-truth trajectory.
5.4 Predictive Check: Individual-Level RMSE
@@
Table 3 presents the predictive check on individual-level RMSE for continuous variables. We aim to assess the performance of G-Transformer, G-Net, and Causal Transformer in predicting 24-hour patient covariate trajectories, starting at time $k$ conditioned on covariates from the previous $k-1$ hours in the ICU. Values reported represent the RMSE of continuous covariates between the predicted and ground-truth trajectories. To account for mismatch in trajectory length due to over- and under-prediction of death and release time, for both predicted and actual trajectories, depending on nature of covariates, post-death timesteps are padded with one of the following values: normalized zero, normalized minimum value in dataset, or normalized logarithmic zero. The post-release timesteps are consistently padded with normalized population mean.
@@
Results in Table 3 indicate that G-Transformer consistently outperforms G-Net and Causal Transformer across different values of $k$. Although Causal Transformer Melnychuk et al. (2022) is not intended to estimate counterfactual outcomes under dynamic treatment regimes, nevertheless, we present the performance from a modified version of Causal Transformer (where treatments depend on past covariate history during simulations) here as a baseline.
@@
Table 3: MIMIC-IV Experiments: Predictive check on continuous variables at individual-level RMSE. $k$ is the simulation start time relative to the ICU admission time. Performance in predicting 24-hour patient covariate trajectories, starting at hour $k$ conditioned on covariates from the previous $k-1$ timesteps of post-ICU admission. Values reported represent the individual-level RMSE of continuous covariates between the predicted and ground-truth trajectories. Individual-level RMSE is adjusted to account for error due to over- and under-prediction in the trajectory length as a result of death and discharge. *Causal Transformer modified to update treatment assignment at each time-step based on predicted action.
Individual Level RMSE
Model / k (hours) 2 6
Causal Transformer 14.37 15.73
G-Net 7.017 7.670
G-Transformer 6.993 7.643
@@
Figure 4: Illustration of predictive check on an example MIMIC test set patient. G-Transformer simulated systolic blood pressure (SBP) and heart rate (HR) trajectories (100 Monte Carlo simulations in light blue, average in solid dark blue) compared with ground truth (red) for one patient with predicted treatments under the observational treatment regime (simulation starting at hour 6).
@@
Table 4 presents AUCs in predicting clinical outcomes within the first 24 hours of ICU admission at varying simulation start times from G-Transformer and G-Net. There are four outcomes of interest: diagnosis of pulmonary edema, use of a mechanical ventilator (MV), administration of diuretics, and dialysis. We observe that the predicted 24-hour clinical outcomes of interest tend to be more accurate as the value of $k$ increases, and G-Transformer performs slightly better than G-Net when conditioned on a longer history (i.e. $k$=6) except in the case of predicting diuretics use.
@@
Table 4: MIMIC-IV Experiments: Predictive check on clinical outcomes within the first 24 hours of ICU admission at varying simulation start times. Values reported are AUCs. Edema represents presence of Pulmonary Edema.
Model k Edema MV Diuretics Dialysis
G-Net 2 0.802 0.898 0.735 0.878
G-Transformer 2 0.817 0.896 0.644 0.880
G-Net 6 0.850 0.938 0.713 0.846
G-Transformer 6 0.864 0.945 0.650 0.958
5.5 Counterfactual Experiments
Our counterfactual strategies are adapted from established clinical trials studying the early treatment of sepsis. This strategy imposes a fluid cap on the total amount of fluids intake of a patient to $X$ liters, and ceases administration once the fluid cap is reached or when the patient is fluid-overloaded. More specifically, this strategy was based on the Crystalloid Liberal or Vasopressors Early Resuscitation in Sepsis (CLOVERS) clinical trial (Self et al., ). For a patient with blood pressure below 65mmHg at time $t$, a 1000mL bolus was administered if the total volume of fluids (including both treatment and maintenance) they received up until that time point did not exceed $X$ liters and if they did not exhibit any signs of fluid overload (as indicated by the presence of pulmonary edema based on chest x-ray radiology reports). We experimented with imposing a fluid cap of $5$L as fluid conservative strategies, and a fluid strategy with no fluid cap to represent a fluid liberal strategy.
Illustrative Examples in Counterfactual Prediction